A card is selected at random from an ordinary deck of 52 playing cards. Let E be the event that the selected card is an ace and F be the event that it is a spade. Are the two events E and F disjoint? Are they independent? Why?
Answer #1

P(E) =-= 52 13 ( i.e. Number of aces are 4)

P(F)=\frac{13}{52}=\frac{1}{4} ( i.e. Number of spades are 13 )

P(E\cap F)=\frac{1}{52}

Therefore E and F are not disjoint

P(E)\times P(F)=\frac{1}{13}\times \frac{1}{4}=\frac{1}{52}=P(E\cap F)

Hence E and F are  independent